OpenStudy (anonymous):

a boat's crew rowed 12 miles downstream with the current in 1.5 hours. the return trip upstream covered the same distance but took 4 hours. find the crews rowing rate in still water and the rate of the current

OpenStudy (rulnick):

r = d / t so ... row + current = 12 / 1.5 = 8 row - current = 12 / 4 = 3 add the two equations to get 2 row = 11 row = 11/2 = 5.5 therefore current = 2.5 (all in miles per hour)

OpenStudy (rulnick):

Does that help? Need more explanation?

OpenStudy (anonymous):

Thank you